1. A solution contains 8.68×10-3 M sodium hydroxide and 6.04×10-3 M ammonium sulfide. Solid manganese(II) nitrate...

1.

A solution contains 8.68×10-3 M sodium hydroxide and 6.04×10-3 M ammonium sulfide.
Solid manganese(II) nitrate is added slowly to this mixture.

A. What is the formula of the substance that precipitates first?

formula =   

B. What is the concentration of manganese(II) ion when this precipitation first begins?
[Mn2+] = M

2.

A solution contains 5.04×10-3 M magnesium nitrate and 1.31×10-2 M nickel(II) acetate.
Solid sodium hydroxide is added slowly to this mixture.

A. What is the formula of the substance that precipitates first?

formula =  

B. What is the concentration of hydroxide ion when this precipitation first begins?
[OH-] = M

Solubility Product Constants (Ksp at 25 oC)
Type   Formula   Ksp
Solubility Product Constants (Ksp at 25 oC)
Type   Formula   Ksp
Bromides   PbBr2   6.3 × 10-6


AgBr   3.3 × 10-13
Carbonates   BaCO3   8.1 × 10-9


CaCO3   3.8 × 10-9


CoCO3   8.0 × 10-13


CuCO3   2.5 × 10-10


FeCO3   3.5 × 10-11


PbCO3   1.5 × 10-13


MgCO3   4.0 × 10-5


MnCO3   1.8 × 10-11


NiCO3   6.6 × 10-9


Ag2CO3   8.1 × 10-12


ZnCO3   1.5 × 10-11
Chlorides   PbCl2   1.7 × 10-5


AgCl   1.8 × 10-10
Chromates   BaCrO4   2.0 × 10-10


CaCrO4   7.1 × 10-4


PbCrO4   1.8 × 10-14


Ag2CrO4   9.0 × 10-12
Cyanides   Ni(CN)2   3.0 × 10-23


AgCN   1.2 × 10-16


Zn(CN)2   8.0 × 10-12
Fluorides   BaF2   1.7 × 10-6


CaF2   3.9 × 10-11


PbF2   3.7 × 10-8


MgF2   6.4 × 10-9
Hydroxides   AgOH   2.0 × 10-8


Al(OH)3   1.9 × 10-33


Ca(OH)2   7.9 × 10-6


Cr(OH)3   6.7 × 10-31


Co(OH)2   2.5 × 10-16


Cu(OH)2   1.6 × 10-19


Fe(OH)2   7.9 × 10-15


Fe(OH)3   6.3 × 10-38


Pb(OH)2   2.8 × 10-16


Mg(OH)2   1.5 × 10-11


Mn(OH)2   4.6 × 10-14


Ni(OH)2   2.8 × 10-16


Zn(OH)2   4.5 × 10-17
Iodides   PbI2   8.7 × 10-9


AgI   1.5 × 10-16
Oxalates   BaC2O4   1.1 × 10-7


CaC2O4   2.3 × 10-9


MgC2O4   8.6 × 10-5
Phosphates   AlPO4   1.3 × 10-20


Ba3(PO4)2   1.3 × 10-29


Ca3(PO4)2   1.0 × 10-25


CrPO4   2.4 × 10-23


Pb3(PO4)2   3.0 × 10-44


Ag3PO4   1.3 × 10-20


Zn3(PO4)2   9.1 × 10-33
Sulfates   BaSO4   1.1 × 10-10


CaSO4   2.4 × 10-5


PbSO4   1.8 × 10-8


Ag2SO4   1.7 × 10-5
Sulfides   CaS   8 × 10-6


CoS   5.9 × 10-21


CuS   7.9 × 10-37


FeS   4.9 × 10-18


Fe2S3   1.4 × 10-88


PbS   3.2 × 10-28


MnS   5.1 × 10-15


NiS   3.0 × 10-21


Ag2S   1.0 × 10-49


ZnS   2.0 × 10-25
Sulfites   BaSO3   8.0 × 10-7


CaSO3   1.3 × 10-8


Ag2SO3   1.5 × 10-14
Answer #1

1)

Ksp of Mn(OH)2 = 4.6 × 10^-14

Ksp of MnS = 5.1 × 10^-15

A.)

formula = MnS

B.)

MnS (s) -------------> Mn2+ (aq) + S2- (aq)

Ksp = [Mn2+][S2-]

5.1 x 10^-15 = [Mn2+] (6.04 x 10^-3)

[Mn2+] = 8.44 x 10^-13 M

2.

Ksp of Mg(OH)2 = 1.5 × 10-11

Ksp of Ni(OH)2 = 2.8 × 10-16

A.)

formula = Ni(OH)2

B.)

Ni(OH)2 (s)   -------------> Ni2+ (aq) + 2 OH- (aq)

Ksp = [Ni2+][OH-]^2

2.8 x 10^-16 = (1.31 x 10^-2) [OH-]^2

[OH-] = 1.46 x 10^-7 M